OpenStudy (anonymous):

Ryan's birthday cake was made by using a cylindrical baking pan . If the top layer has a 10 inch diameter and is 6 inches tall , what is the capacity of the pan , to the nearest cubic inch ?

OpenStudy (anonymous):

rayon: 5in volume: pi*r^2*h =pi*5^2*6 =471cubic inch

OpenStudy (anonymous):

thanks . and hey whats 2,010.619298297468 to the nearest cubin inch ? @d91196

OpenStudy (anonymous):

2,011